Construct a scenario where Barium fluoride is precipitating.

Barium fluoride (BaF2) Ksp=1.7x10^-6
Answer #1

Assuming we have a liter of solution, to which we add M moles of BaF2:

[BaF2] = moles / Volume = M / 1 = M

[BaF2] = [Ba2+] = M

[F-] = 2 * [BaF2] = 2M

Now using the expression for Ksp= [Ba2+] * [F-] 2

1.7 * 10-6 = M * (2M)2

1.7 * 10-6 = 4M3

Solving, M = 7.52 * 10-3 moles

Mass of BaF2 added = moles * molar mass = 7.52 * 10-3 * 175.3 = 1.32 g

Therefore we add 1.32 g of this salt to a liter of water to make it just precipitate